### 16 Mar 2026 v1

# A deep backward regression-based scheme for high-dimensional nonlinear partial differential equations

Abstract. A deep backward regression-based (DBR) scheme for solving high-dimensional nonlinear parabolic partial differential equations is proposed. Building upon the established DBDP method (Huré et al., 2020), our algorithm introduces a reformulation of the local loss functions that are sequentially optimized via backward induction at each time step. The core of this approach involves reformulating simulated backward stochastic difference equations into their conditional expectation representations, thereby recasting a projection-based stochastic optimization problem as a deterministic function-approximation task. By explicitly incorporating conditional expectations, the DBR scheme facilitates a denoising mechanism prior to loss evaluation. This architecture substantially mitigates numerical variance, resulting in enhanced training stability and superior generalization performance. Numerical results demonstrate that the DBR scheme consistently outperforms the DBDP1 method, maintaining accuracy up to $d=200$ for bounded solutions (see Table 1). Notably, for complex unbounded PDEs where the DBDP1 method fails beyond $d=10$, the DBR scheme remains robust up to $d=20$ with relative errors under $9.7\%$ (see Table 6). Theoretically, we derive rigorous upper error bounds and establish half-order convergence for the proposed scheme. Extensions to variational inequalities are also provided.
Key words. high dimensional nonlinear PDEs, backward stochastic differential equations, conditional expectations, deep neural networks, numerical approximations
AMS subject classifications. 60H35, 65C20, 65M12
1 Introduction
In this paper, we develop a class of artificial neural networks based on probabilistic numerical methods to approximate the solutions of high-dimensional quasilinear parabolic PDEs of the following form
[…]
where $\mu$ is a map from $0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $\sigma$ is a map from $0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$; $f:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ represent the nonlinear generator and the terminal function respectively; the operators $D_{x}$ and $D_{x}^{2}$ refer to the first and second order spatial derivatives; the symbol $\cdot$ represents the scalar product and denotes the transpose of vector or matrix.
The pioneer works of Pardoux and Peng [31, 32], Peng [33] show that the quasilinear parabolic PDEs (1.1) are associated with Markovian forward backward stochastic differential equations (FBSDEs) due to the nonlinear Feynman-Kac formula. Since then, significant interest has been directed toward the probabilistic interpretation of PDE solutions. By these interpretations, researchers have developed probabilistic numerical methods to approximate PDE solutions (see [9, 10, 12] and references therein). Concurrently, the design of algorithms to solve FBSDEs has been intensively studied. Up to now, the proposed numerical algorithms to solve nonlinear FBSDEs have shown their remarkable performance with respect to the high-order problems. For instance, the Euler schemes with convergence rates of $\frac{1}{2}$ or $1$ are presented in [2, 19, 36, 37]. Multi-step schemes [24, 25, 21, 38] achieve convergence rates greater than or equal to $1$, although they require terminal conditions over multiple steps. Furthermore, one-step schemes with convergence rates of at least $2$ have been proposed in [8, 35, 26].
However, most of these numerical methods are unsuitable for solving nonlinear PDEs in dimensions greater than 4 (see [4]). To address this issue, some techniques such as the parallel computing method, the sparse grid method, and the variance reduction method(including the control variate methods, the importance sampling methods, and the multilevel Monte Carlo methods et al.) have proven effective for nonlinear PDEs with the number of dimension below 20 (see [20]). The high dimensional problem (that is the so-called curse of dimensionality) remains a major challenge in the numerical solutions of PDEs, even in the broader field of scientific computing.
Until 2017, E, Han and Jentzen [11, 22] introduce the first deep learning based on numerical method, called the deep BSDE method which uses the required values as parameters and utilizes the feedforward neural networks to approximate the solutions through solving an optimization problem and demonstrate the effectiveness in 100-dimensional settings by conducting various numerical experiments. To date, there are lots of numerical algorithms based on neural networks. They are designed by adjusting, reformulating or extending the deep BSDE method to solve high dimensional PDEs (see [13, 17, 28, 23, 29, 34, 15, 16] and many others) and these deep numerical algorithms have also shown their remarkable performance on the high-dimensional problems. Recently, Cai, Fang and Zhou [5, 7] introduce the SOC-MartNet method to solve ultra-high-dimensional quasilinear parabolic PDEs and it demonstrates the strong performance on large-scale benchmarks. Furthermore, they propose the deep random difference method in [6] to reduce variance and improve stability of the SOC-MartNet method.
Among the literature on deep learning algorithms for high-dimensional PDEs, an important DBDP method was proposed by Hure, Pham and Warin in [28]. It estimates the solutions and their gradient by the loss functions, for $i=N-1,N-2,\cdots,0$
$J_{i}^{HPW}(\theta_{i})=\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{Y}_{i+1}^{*}-hf(t_{i},X_{i}^{\pi},\mathcal{Y}_{i},\mathcal{Z}_{i})+\mathcal{Z}_{i}\Delta W_{i}\right|^{2}\right],$ (1.2)
at each time step $t_{i}$, where $\mathcal{Y}_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ and $\mathcal{Z}_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n\times d}$ denote the sequences of network functions. Then update $(\mathcal{Y}_{i}^{*},\mathcal{Z}_{i}^{*})$ as the solutions of the local minizamition problems. Thus, we only need to consider the local gradient as updating the parameter of neural networks, which significantly reduces the difficulty of backpropagation and solves the training difficulties brought by long-sequence time dependencies. The loss function (1.2) is based on the Euler discretization of the BSDE. It measures the residual of the discretized Euler scheme which implies that during the training process, one can directly assess the accuracy of the scheme by observing whether the loss value approaches zero. But the random noise term $\mathcal{Z}_{i}\Delta W_{i}$ in (1.2) may lead to fluctuations especially when $\mathcal{Z}_{i}$ is large (such as high volatility or complex hedging requirements). Naturally, it may cause oscillations in the gradients which lead to divergence. Besides, the loss function $J_{i}^{HPW}(\theta_{i})$ is prone to overfitting the noise of the simulated paths, leading to poor generalization ability.
To address these problems, we propose a new deep backward regression-based (DBR) method to approximate the solutions and their gradient of high-dimensional PDEs by the loss functions
$F_{y,i}(\theta_{y,i})=\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathbb{E}_{i}\left[\mathcal{Y}_{i+1}^{*}+hf(t_{i},X_{i}^{\pi},\mathcal{Y}_{i},\mathcal{Z}_{i}^{*})\right]\right|^{2}\right],$ (1.3)
and
$F_{z,i}(\theta_{z,i})=\mathbb{E}\left[\left|\mathcal{Z}_{i}-\mathbb{E}_{i}\left[\mathcal{Y}_{i+1}^{*}\frac{\Delta W_{i}^{\top}}{h}\right]\right|^{2}\right],$ (1.4)
where $\mathbb{E}_{i}\left[\cdot\right]=\mathbb{E}\left[\cdot\big|\mathcal{F}_{t_{i}}\right]$ is the conditional expectation. The above regression-based loss functions which utilize conditional expectations are different from those proposed in Remark 2.3 of Germain, Pham and Warin [17]. The accuracy of the DBR scheme can also be tested when computing at each time step the infimum of loss function, which should be equal to zero as the DBDP scheme in [17]. By explicitly incorporating the conditional expectations, the DBR scheme effectively performs a ‘denoising’ step prior to loss evaluation. In terms of implementation, this is typically realized through multi-path Monte Carlo averaging. Its essence lies in replacing the simulated backward stochastic difference equations with their conditional expectation forms and transforming a projection-based stochastic optimization problem into a smoother deterministic function-fitting task.
To achieve more effective approximations of the solutions of PDEs, we design the DBR algorithm via the following manners: (i) the FBSDEs is discretized by the Euler scheme which is introduced by Zhang [36]; (ii) utilize the artificial neural networks equipped with the input variables $(t_{i},X_{i}^{\pi})$ to learn the conditional expectations at each time step $t_{i}$; (iii) adopt a step-by-step approach, optimizing the variables $Z_{i}$ first and then the variables $Y_{i}$ at each time step, rather than simultaneously optimizing the pairs $(Y_{i},Z_{i})$ ; (iv) introduce the loss functions (1.3) and (1.4) to train the deep feedforward neural networks.
Moreover, the core of the DBR method is to approximate the conditional expectation through multi-path Monte Carlo averaging. This design shows some advantages when dealing with high-dimensional or complex-structured PDEs: (i) noise smoothing: multi-path averaging mitigates the impact of single-path Brownian noise $\Delta W_{i}$, reducing numerical fluctuations (e.g., the loss function of regression-based methods does not directly depend on $\Delta W_{i}$). (ii) generalization ability: the DBR method learns “local solutions of multi-paths” rather than “local solutions of single paths”, making them more likely to avoid overfitting. Especially, as the number of dimensions of PDEs with an unbounded solution and complex structure exceeds 10, the DBR method effectively mitigates the adverse effects of increasing dimensionality (see Tables 5 and 6). The method’s consistent performance at dimensions $d=15$ and $d=20$, particularly when contrasted with the breakdown of DBDP1, provides compelling evidence of its practical utility for high-dimensional problems.
The remainder of this paper is organized as follows. In Section 2, some essential materials that are used in the whole paper are provided. The DBR method is designed in Subsection 3.1; The upper bound error of the proposed DBR method is rigorously derived theoretically in Subsection 3.2. In addition, we provide the convergence analysis for the proposed DBR method in Subsection 3.3. In Section 4, we extend the DBR scheme to variational inequalities and prove its corresponding convergence. We report the numerical results in Section 5. In the end, we compile the conclusion of this paper in Section 6.
2 Preliminaries
In this section, we present some materials that will be utilized to design the deep probabilistic numerical approximation method for the high-dimensional nonlinear PDEs, such as the probabilistic representation of the solutions to the PDEs, the function approximations in the class of neural networks and the Euler scheme for FBSDEs.
2.1 Deep neural networks as function approximators
The universal approximation theorem in [27] indicates that we can use the neural networks to approximate the unknown functions and this is also reasonable. Therefore, we recall some notations and basic definitions with respect to the feedforward neural networks which are a type of deep neural networks and will be applied in the following parts. Let
$\mathcal{LL}_{d_{1},d_{2}}^{\varrho}=\left\{\psi:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}^{d_{2}}:\exists~(\mathcal{W},\beta)\in\mathbb{R}^{d_{2}\times d_{1}}\times\mathbb{R}^{d_{2}},\psi(x)=\varrho(\mathcal{W}x+\beta)\right\}$
represent the set of the layer functions with the input dimension $d_{1}$, the output dimension $d_{2}$ and the nonlinear function (called the activation function) $\varrho:\mathbb{R}^{d_{2}}\rightarrow\mathbb{R}^{d_{2}}$. The activation function $\varrho$ works by utilizing componentwise the one dimensional activation function, namely $\varrho(x_{1},x_{2},\ldots,x_{d_{2}})=(\overline{\varrho}(x_{1}),\overline{\varrho}(x_{2}),\ldots,\overline{\varrho}(x_{d_{2}}))$, with $\overline{\varrho}(x):\mathbb{R}\rightarrow\mathbb{R}$, to the affine map $x\in\mathbb{R}^{d_{1}}\rightarrow\mathcal{W}x+\beta\in\mathbb{R}^{d_{2}}$ with the weight $\mathcal{W}\in\mathbb{R}^{d_{2}\times d_{1}}$ and the bias $\beta\in\mathbb{R}^{d_{2}}$. We write $\mathcal{LL}_{d_{1},d_{2}}^{\varrho}$ as $\mathcal{LL}_{d_{1},d_{2}}$ when $\varrho$ denotes the identity function. Thus, the set of the feedforward neural networks can be defined as
$\mathcal{NN}_{\mathbf{d}_{0},\mathbf{d}}^{\varrho,L,\mathbf{n}}=$ $\left\{\varphi:\mathbb{R}^{\mathbf{d}_{0}}\rightarrow\mathbb{R}^{\mathbf{d}}:\exists~\psi_{0}\in\mathcal{LL}_{\mathbf{d}_{0},n_{0}}^{\varrho_{0}},\exists~\psi_{\ell}\in\mathcal{LL}_{n_{\ell-1},n_{\ell}}^{\varrho_{\ell}},\ell=1,2\ldots,L-1,\right$.
$\left.\exists~\psi_{L}\in\mathcal{LL}_{n_{L-1},\mathbf{d}},\varphi=\psi_{L}\circ\psi_{L-1}\circ\cdots\circ\psi_{0}\right\}$
with the input dimension $\mathbf{d}_{0}$, the output dimension $\mathbf{d}$ and $L$ hidden layers with $n_{\ell}$ neurons in the $\ell$-th layer for $\ell=0,1,\ldots,L-1$. Naturally, the coefficients $\mathbf{d}_{0},\mathbf{d},L,\mathbf{n}=\{n_{\ell}\}_{0\leq\ell\leq L-1}$ and the activation $\varrho=\{\varrho_{\ell}\}_{0\leq\ell\leq L-1}$ form the architecture of the network.
2.2 Probabilistic representation for the solutions of PDEs
Our deep probabilistic numerical method for the PDEs (1.1) relies on FBSDEs representation of its solutions. The well-known nonlinear Feynman-Kac formula (see [31, 33, 32]) shows that on the filtered complete probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{P})$, the solution $(Y,Z)\in\mathbb{R}^{n}\times\mathbb{R}^{n\times d}$ of the FBSDEs
[…]
related to the solution $u$ of the PDEs (1.1) via
$Y_{t}=u(t,X_{t}),\qquad Z_{t}=\sigma^{\top}(t,X_{t})D_{x}u(t,X_{t}),$ (2.2)
where $u(s,x)\in C_{b}^{1,2}([0,T]\times\mathbb{R}^{d})$; $(\mathcal{F}_{t})_{0\leq t\leq T}$ denotes the standard Brownian filtration; $W_{t}$ denotes a $d$-dimensional Brownian motion; $x_{0}\in\mathbb{R}^{d}$ represents a given initial value of the stochastic differential equation (SDE) in (2.1).
For readers’ convenience, before giving the time-discretization scheme of FBSDEs (2.1), we introduce the following notations. We denote the grid of the time interval $[0,T]$ by $\pi$, namely $\pi=\{0=t_{0}<t_{1}<\cdots<t_{N}=T\}$, where $t_{i}=ih,h=\frac{T}{N}$ for $i=0,1,\cdots,N,N\in\mathbb{N}^{+}$ and $\Delta W_{i}=W_{t_{i+1}}-W_{t_{i}}$ the $(i+1)$-th Brownian motion increment. Then the Euler time-discretization of the the backward stochastic differential equation (BSDE) in (2.1), at the mesh points $t_{i}$, is as follows
$Y_{i}^{\pi}=Y_{i+1}^{\pi}+hf(t_{i},X_{i}^{\pi},Y_{i}^{\pi},Z_{i}^{\pi})-Z_{i}^{\pi}\Delta W_{i},$ (2.3)
which also reads as the conditional expectation formula (see [36])
$\left\{\begin{aligned} Y_{i}^{\pi}=&\mathbb{E}_{i}\Big[Y_{i+1}^{\pi}+hf(t_{i},X_{i}^{\pi},Y_{i}^{\pi},Z_{i}^{\pi})\Big],\\ Z_{i}^{\pi}=&\mathbb{E}_{i}\big[Y_{i+1}^{\pi}\frac{\Delta W_{i}^{\top}}{h}\big],\end{aligned}\right$. (2.4)
where $X_{i+1}^{\pi}=X_{i}^{\pi}+h\mu(t_{i},X_{i}^{\pi})+\sigma(t_{i},X_{i}^{\pi})\Delta W_{i}$; $X_{i}^{\pi}$, $Y_{i}^{\pi}$ and $Z_{i}^{\pi}$ denote the time-discretization form of $X$, $Y$ and $Z$ at $t_{i}$ respectively.
3 Deep backward regression-based method
In this section, we formulate the DBR method to approximate the conditional expectations in scheme (2.4) using learned data. Our aim is to provide a fully implementable algorithm. Subsequently, we present a comprehensive error analysis of the DBR method. Finally, we investigate its convergence rate.
3.1 Deep learning method to approximate the conditional expectation
To implement $(Y_{i}^{\pi},Z_{i}^{\pi})$ based on the temporal discretization scheme (2.4), one has to approximate the conditional mathematics expectations. Now, we obtain them by means of the deep neural network method.
In what follows, the forward process $X_{\cdot}^{\pi,m}$ is simulated by Monte Carlo method in the following manners, for $i=0,1,\cdots,N-1$
$X_{i+1}^{\pi,m}=X_{i}^{\pi,m}+\mu(t_{i},X_{i}^{\pi,m})h+\sigma(t_{i},X_{i}^{\pi,m})\Delta W_{i}^{m},\quad X_{0}^{\pi,m}=x_{0},$ (3.1)
where $\Delta W_{i}^{m}\sim\mathcal{N}(0,h)$, $m=1,2,\cdots,M$, $M\in\mathbb{N}$. The backward process $Y_{\cdot}^{\pi,m}$ is simulated by Monte Carlo method in the following ways, for $i=N-1,N-2,\cdots,0$
$Y_{i}^{\pi,m}=Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})-Z_{i}^{\pi,m}\Delta W_{i}^{m},$ (3.2)
which also expresses as the conditional expectation form
$\left\{\begin{aligned} Y_{i}^{\pi,m}=&\mathbb{E}_{i}\Big[Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})\Big],\\ Z_{i}^{\pi,m}=&\mathbb{E}_{i}\big[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\big].\end{aligned}\right$. (3.3)
At each time step $t_{i}$, the forward process $X_{i}^{\pi,m,k}$ is generated as below:
$X_{i+1}^{\pi,m,k}=X_{i}^{\pi,m}+\mu(t_{i},X_{i}^{\pi,m})h+\sigma(t_{i},X_{i}^{\pi,m})\Delta W_{i}^{m,k},$
where $\Delta W_{i}^{m,k}\sim\mathcal{N}(0,h)$ is independently and identically distributed as $\Delta W_{i}^{m}$; $k=1,2,\cdots,K_{i}$; $K_{i}\in\mathbb{N}$ denotes the number of trajectories in the Monte Carlo simulation.
Now, the approximation algorithm is implemented in detail. The terms on the right hand side of (3.3) are written as
[…]
by Monte Carlo method, where $Y_{i+1}^{\pi,m,k}=u(t_{i+1},X_{i+1}^{\pi,m,k})$ is independently and identically distributed as $Y_{i+1}^{\pi,m}$;
$R_{i,1}^{y}=$ $\mathbb{E}_{i}\Big[Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})\Big]-\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}-hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m}),$
$R_{i,1}^{z}=$ $\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]-\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right].$
Replacing the two random variables in local Monte Carlo approximates in (3.4) with two deep feedforward neural networks, namely
[…]
where $p_{i}^{m}=(t_{i},X_{i}^{\pi,m})$; $\theta^{0}_{y,i}$ and $\theta^{0}_{z,i}$ denote the initial parameters of the two neural networks at each time step;
$R_{i,2}^{y,0}=$ $\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{0}_{y,i}),$
$R_{i,2}^{z,0}=$ $\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right]-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{0}_{z,i}).$
From (3.4) and (3.5), we have
$\left\{\begin{aligned} Y_{i}^{\pi,m}=&\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{0}_{y,i})+R_{i,2}^{y,0}+R_{i,1}^{y},\\ Z_{i}^{\pi,m}=&\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{0}_{z,i})+R_{i,2}^{z,0}+R_{i,1}^{z}.\end{aligned}\right$. (3.6)
To train the deep feedforward neural networks, we introduce the expectation loss functions in the following forms
$\left\{\begin{aligned} F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})=&\frac{1}{M}\sum_{m=1}^{M}\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2},\\ F_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})=&\frac{1}{M}\sum_{m=1}^{M}\left|S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2},\end{aligned}\right$. (3.7)
where $p_{i}^{m,k}=(t_{i},X_{i}^{\pi,m,k})$; $r_{j}\in\mathbb{N}^{+}$ denotes the number of iterations of the stochastic gradient descent (SGD) method; $\theta^{r_{j}}_{y,i}=\theta_{y,i}^{r_{j}-1}-\rho_{y,i}^{r_{j}-1}\nabla_{\theta_{y,i}^{r_{j}-1}}F_{y,i}^{r_{j}-1}(\theta_{y,i}^{r_{j}-1}),$ $\theta^{r_{j}}_{z,i}=\theta_{z,i}^{r_{j}-1}-\rho_{z,i}^{r_{j}-1}\nabla_{\theta_{z,i}^{r_{j}-1}}F_{z,i}^{r_{j}-1}(\theta_{z,i}^{r_{j}-1})$; $\rho_{y,i}^{r_{j}},\rho_{z,i}^{r_{j}}\in(0,+\infty)$ denote the learning rates; $\theta^{*}_{y,i}=\arg\min\limits_{\theta_{y,i}^{r_{j}}}F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})$, $\theta^{*}_{z,i}=\arg\min\limits_{\theta_{z,i}^{r_{j}}}F_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})$;
$S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m,k},\theta_{y,i+1}^{*})\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right],$
$S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m,k},\theta_{y,i+1}^{*})$
$\quad+hf(t_{i},X_{i}^{\pi,m},\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}}),\mathcal{NN}_{z,i+1,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})).$
Thus, by the optimization of the SGD method, the (3.6) is rewritten as
$\left\{\begin{aligned} Y_{i}^{\pi,m}=&\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i})+R_{i,2}^{y}+R_{i,1}^{y},\\ Z_{i}^{\pi,m}=&\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{z,i})+R_{i,2}^{z}+R_{i,1}^{z},\end{aligned}\right$. (3.8)
with
$R_{i,2}^{y}$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{*}),$
$R_{i,2}^{z}$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right]-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*}).$
Therefore, we deduce the fully discrete approximation solutions $(\mathcal{Y}_{i},\mathcal{Z}_{i})$ for the analytical solutions $(Y_{t_{i}},Z_{t_{i}})$ for $i=N,N-1,\cdots,1,0$, $m=1,2,\cdots,1,M$ based on the DBR method as below
  1. 1.
    the terminal condition is $(\mathcal{Y}_{N},\mathcal{Z}_{N})=(\Phi(X_{N}^{\pi,m}),\sigma^{\top}(t_{N},X^{\pi,m}_{N})D_{x}u(t_{N},X^{\pi,m}_{N}))$,
  2. 2.
    for $0\leq i<N,$ the transition from $i+1$ to $i$ is given by
    $\left\{\begin{aligned} \mathcal{Y}_{i}=&\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i}),\\ \mathcal{Z}_{i}=&\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{z,i}).\end{aligned}\right$. (3.9)
In summary, the DBR algorithm is provided to compute the numerical solutions $\mathcal{Y}_{i}$ and $\mathcal{Z}_{i}$ (see Algorithm 1).
Input: Grid $0=t_{0}<\dots<t_{N}=T$ with step $h=T/N$.
Output: Trained networks $\{\mathcal{NN}_{y,i},\mathcal{NN}_{z,i}\}_{i=0}^{N-1}$ and learnable parameters.
Initialize: Networks $\theta_{y,i},\theta_{z,i}\sim$ Xavier initialization [18];
for $i=N-1,\dots,0$ do
 for $m=1,\dots,M$ do
      Generate $\Delta W_{i}^{m}\sim\mathcal{N}(0,h)$;
    $X^{\pi,m}_{i+1}\leftarrow X_{i}^{\pi,m}+b(t_{i},X_{i}^{\pi,m})h+\sigma(t_{i},X_{i}^{\pi,m})\Delta W_{i}^{m}$, $X_{0}=x_{0}$;
   end for
  Generate $\Delta W_{i}^{m,k}\sim\mathcal{N}(0,h)$, for $m\in\{1,\cdots,M\},k\in\{1,\cdots,K\}$;
   Compute next states: $X_{i+1}^{\pi,m,k}\leftarrow X^{\pi,m}_{i}+b(t_{i},X^{\pi,m}_{i})h+\sigma(t_{i},X^{\pi,m}_{i})\Delta W_{i}^{m,k}$;
 if i==N1i==N-1 then
    $\mathcal{Y}_{i+1}^{m,k}\leftarrow g(t_{i+1},X_{i+1}^{\pi,m,k})$;
 else
    $\mathcal{Y}_{i+1}^{m,k}\leftarrow\mathcal{NN}_{y,i+1}(t_{i+1},X_{i+1}^{\pi,m,k},\theta^{*}_{y,i})$;
   end if
 $Loss_{Z,i}\leftarrow\frac{1}{M}\sum\limits_{m=1}^{M}\left[\left|\frac{1}{K}\sum\limits_{k=1}^{K}\left(\mathcal{Y}_{i+1}^{m,k}\frac{\Delta W_{i}^{m,k}}{h}\right)-\mathcal{NN}_{z,i}(t_{i},X_{i}^{\pi,m},\theta_{z,i})\right|^{2}\right]$;
 Update $\theta_{z,i}$ using the SGD method and the optimal parameter is denoted as $\theta^{*}_{z,i}$;
 $\mathcal{Z}_{i}^{m,k}\leftarrow\mathcal{NN}_{z,i}(t_{i},X_{i}^{\pi,m,k},\theta_{z,i}^{*})$;
 $\mathcal{Z}_{i}^{m}\leftarrow\mathcal{NN}_{z,i}(t_{i},X_{i}^{\pi,m},\theta_{z,i}^{*})$;
 $Loss_{Y,i}\leftarrow\frac{1}{M}\sum\limits_{m=1}^{M}\Bigg[\bigg|\frac{1}{K}\sum\limits_{k=1}^{K}\mathcal{Y}_{i+1}^{m,k}+f(t_{i},X_{i}^{\pi,m,k},\mathcal{NN}_{y,i}(t_{i},X_{i}^{\pi,m},\theta_{y,i}^{r}),\mathcal{Z}_{i}^{m,k})\cdot h$ $\mathcal{NN}_{y,i}(t_{i},X_{i}^{\pi,m},\theta_{y,i})\bigg|^{2}\Bigg]$;
 Update $\theta_{y,i}$ using the SGD method and the optimal parameter is denoted as $\theta^{*}_{y,i}$;
 $\mathcal{Y}_{i}^{m}\leftarrow\mathcal{NN}_{y,i}(t_{i},X_{i}^{\pi,m},\theta_{y,i}^{*})$;
end for
Final Estimation: $\mathcal{Y}_{0}=\mathcal{NN}_{y,0}(t_{0},X_{0},\theta^{*}_{y,0})$;
Algorithm 1 DBR method for FBSDEs
The DBDP method in [28] uses $\mathcal{Y}_{i+1}^{*}+hf(t_{i},X_{i}^{\pi},\mathcal{Y}_{i},\mathcal{Z}_{i})-\mathcal{Z}_{i}\Delta W_{i}$ as the label for $(\mathcal{Y}_{i}(\theta_{i}),\mathcal{Z}_{i}(\theta_{i}))$ ( or $(\mathcal{Y}_{i}(\theta_{i}),\sigma^{\top}(t_{i},\cdot)D_{x}\mathcal{Y}_{i}(\theta_{i})$) at each time step $t_{i}$. $\mathcal{Y}_{i}$ and $\mathcal{Z}_{i}$ are $\mathcal{F}_{t_{i}}$ measurable while the information of Brownian motion at moment $t_{i+1}$ is contained in the label. This causes some fluctuations in the estimates of $\mathcal{Y}_{i}$ and $\mathcal{Z}_{i}$. Our DBR method employs $\mathbb{E}_{i}\left[\mathcal{Y}_{i+1}^{*}\right]+hf(t_{i},X_{i}^{\pi},\mathcal{Y}_{i},\mathcal{Z}_{i}^{*})$ and $\mathbb{E}_{i}\left[\mathcal{Y}_{i+1}^{*}\frac{\Delta W_{i}^{\top}}{h}\right]$ as the labels for $\mathcal{Y}_{i}$ and $\mathcal{Z}_{i}$ respectively. This can smooth the noise $\Delta W_{i}$ by averaging, and then makes the estimation $\mathcal{Y}_{i}$ and $\mathcal{Z}_{i}$ more stable at each time step. The DBDP method relies on the information of single-path for labeling, which is random and prone to learning the incorrect solutions of PDEs. Our DBR method, by multi-path averaging, offsets the noise of single-path. This makes our method more robust and having stronger generalization ability (especially in high-dimensional or complex PDEs).
3.2 Error analysis
We now estimate the error of the numerical solutions $(\mathcal{Y}_{i},\mathcal{Z}_{i})_{0\leq i\leq N-1}$ and the analytical solutions $(Y_{t_{i}},Z_{t_{i}})_{0\leq i\leq N-1}$. This result provides an error estimate of the DBR method (3.9) in terms of the $L^{2}$-sense.
The standard assumptions which guarantee the existence and uniqueness on the coefficients of the FBSDEs (2.1) associated to the semilinear PDEs (1.1) are made as following:
$x_{0}$ is square integrable.
$\mu$ and $\sigma$ are Lipschitz continuous of the spatial variable $x$; assume
$\sup\limits_{0\leq t\leq T}\{|\mu(t,0)|+|\sigma(t,0)|\}\leq L,$
where $L>0$ denotes the Lipschitz constant.
$f$ is $\frac{1}{2}$-Hölder continuous with respect to the time variable $t$ and uniformly Lipschitz continuous in all other variables
$|f(t_{1},x_{1},y_{1},z_{1})-f(t_{2},x_{2},y_{2},z_{2})|\leq L(|t_{1}-t_{2}|^{\frac{1}{2}}+|x_{1}-x_{2}|+|y_{1}-y_{2}|+|z_{1}-z_{2}|),$
for all $t\in[0,T],$ $x_{1},x_{2}\in\mathbb{R}^{d}$, $y_{1},y_{2}\in\in\mathbb{R}^{n}$ and $z_{1},z_{2}\in\mathbb{R}^{n\times d}$; and suppose
$\sup\limits_{0\leq t\leq T}\{|f(t,0,0,0)|\}\leq L.$
$g$ is a linear growth function.
Now, we investigate the errors of the scheme (3.9) and define, for $i=N-1,N-2,\cdots,0$
$\left\{\begin{aligned} \mathcal{U}_{t_{i}}^{m}&:=\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})+hf\left(t_{i},X_{i}^{\pi,m},\mathcal{U}_{t_{i}}^{m},\mathcal{V}_{t_{i}}^{m}\right)\right],\\ \mathcal{V}_{t_{i}}^{m}&:=\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right].\end{aligned}\right$. (3.10)
Theorem 3.1
Suppose the assumptions (i)-(iv) hold. Let $(Y_{t_{i}},Z_{t_{i}})$ and $(\mathcal{Y}_{i},\mathcal{Z}_{i})$ be solutions of the FBSDEs (2.1) and solutions of the DBR method (3.9) respectively. Then, we have, for small enough $h$
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]+\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]$
$\leq C\mathbb{E}\left[\left|g(X_{T})-g(X_{N}^{\pi,m})\right|^{2}\right]+Ch+C\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]$
$\quad+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|R_{i+1,2}^{y}\right|^{2}+\left|R_{i,2}^{z}\right|^{2}+N\left|R_{i+1,1}^{y}\right|^{2}+\left|R_{i,1}^{z}\right|^{2}\right],$
where $C$ represents a positive generic constant which is independent of $\pi$ and may change from line to line.
Proof. Step 1. On the uniform discrete mesh $\pi$, we can write the BSDE in (2.1) at the mesh points $t_{i}$ as follows
$Y_{t_{i}}=Y_{t_{i+1}}+\int_{t_{i}}^{t_{i+1}}f_{s}ds-\int_{t_{i}}^{t_{i+1}}Z_{s}dW_{s},$ (3.11)
where $f_{s}=f(s,X_{s},Y_{s},Z_{s})$. From (3.10) and (3.11), we have, for $i\in\{0,1,\ldots,N-1\}$
$Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}=$ $\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]$ (3.12)
$+\mathbb{E}_{i}\left[\int_{t_{i}}^{t_{i+1}}\left(f_{s}-f\left(t_{i},X_{i}^{\pi,m},\mathcal{U}^{m}_{t_{i}},\mathcal{V}_{t_{i}}^{m}\right)\right)ds\right].$
From Young inequality in the form: $(a+b)^{2}\leq(1+\gamma h)a^{2}+(1+\frac{1}{\gamma h})b^{2},$ for $a,b\in\mathbb{R}$, and some $\gamma>0$ which will be given later, Cauchy-Schwarz inequality, and the Lipschitz condition with respect to $f$, we obtain
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]$
$\leq(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+(1+\frac{1}{\gamma h})\mathbb{E}\left[\left|\mathbb{E}_{i}\left[\int_{t_{i}}^{t_{i+1}}\left(f_{s}-f\left(t_{i},X_{i}^{\pi,m},\mathcal{U}^{m}_{t_{i}},\mathcal{V}^{m}_{t_{i}}\right)\right)ds\right]\right|^{2}\right]$
$\leq(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+h(1+\frac{1}{\gamma h})\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}\left(f_{s}-f\left(t_{i},X_{i}^{\pi,m},\mathcal{U}^{m}_{t_{i}},\mathcal{V}^{m}_{t_{i}}\right)\right)^{2}ds\right]$
$\leq(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+4L^{2}h(1+\frac{1}{\gamma h})\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}\left(|s-t_{i}|+|X_{s}-X_{i}^{\pi,m}|^{2}+|Y_{s}-\mathcal{U}^{m}_{t_{i}}|^{2}+|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}\right)ds\right].$ (3.13)
Because the functions $\mu$ and $\sigma$ are the Lipschitz continuity (see the assumptions (i)), we have (see [30])
$\max_{0\leq i\leq N-1}\mathbb{E}\left[\sup_{s\in\left[t_{i},t_{i+1}\right]}\left|X_{s}-X_{i}^{\pi,m}\right|^{2}\right]\leq Ch,$ (3.14)
$\mathbb{E}\left[\sup_{0\leq t\leq T}\left|X_{t}\right|^{p}\right]\leq C_{p}(1+|x_{0}|^{p}),\quad p\geq 1,$ (3.15)
where $C_{p}$ is a constant depending on $p,\mu,\sigma$ and $T$. From the assumptions (ii)-(iii) and (3.15), we have (see [19, 36])
$\max_{0\leq i\leq N-1}\mathbb{E}\left[\sup_{t\in\left[t_{i},t_{i+1}\right]}\left|Y_{t}-Y_{t_{i}}\right|^{2}\right]$ $\leq Ch,$ (3.16)
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]$ $=\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+h\mathbb{E}\left[|\bar{Z}_{t_{i}}-\mathcal{V}^{m}_{t_{i}}|^{2}\right],$ (3.17)
$\mathbb{E}\left[\int_{0}^{T}f_{s}^{2}ds\right]$ $\leq+\infty,$ (3.18)
where $\bar{Z}_{t_{i}}=\mathbb{E}_{i}\left[\frac{1}{h}\int_{t_{i}}^{t_{i+1}}Z_{t}\mathrm{~d}t\right]$. From (3.14), (3.16) and (3.17), we can rewrite (3.13) as
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\leq$ $(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$ (3.19)
$+4L^{2}h(1+\frac{1}{\gamma h})\left(Ch^{2}+2h\mathbb{E}\left[|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}|^{2}\right]\right)$
$+4L^{2}h(1+\frac{1}{\gamma h})\left(\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+h\mathbb{E}\left[|\bar{Z}_{t_{i}}-\mathcal{V}^{m}_{t_{i}}|^{2}\right]\right).$
Multiplying (3.11) with $\Delta W_{i}^{\top}$ and then taking the conditional expectation on both sides of the derived equation, we obtain
$0=\mathbb{E}_{i}[Y_{t_{i+1}}\Delta W_{i}^{\top}]+\mathbb{E}_{i}\left[\Delta W_{i}^{\top}\int_{t_{i}}^{t_{i+1}}f_{s}ds\right]-\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{i}[Z_{s}]ds.$ (3.20)
From (3.10) and (3.20), we have, for $i\in\{0,1,\ldots,N-1\}$
$\bar{Z}_{t_{i}}-\mathcal{V}^{m}_{t_{i}}=$ $\mathbb{E}_{i}\left[\left(Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right)\frac{\Delta W_{i}^{\top}}{h}\right]+\mathbb{E}_{i}\left[\frac{\Delta W_{i}^{\top}}{h}\int_{t_{i}}^{t_{i+1}}f_{s}ds\right]$
$=$ $\mathbb{E}_{i}\left[\left(Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})-\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right)\frac{\Delta W_{i}^{\top}}{h}\right]$
$+\mathbb{E}_{i}\left[\frac{\Delta W_{i}^{\top}}{h}\int_{t_{i}}^{t_{i+1}}f_{s}ds\right].$
From Cauchy-Schwarz inequality and the tower property of conditional expectations, one yields
$\mathbb{E}\left[\left|\bar{Z}_{t_{i}}-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right]\leq$ $\frac{2d}{h}\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]$ (3.21)
$-\frac{2d}{h}\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]+2d\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right].$
Plugging (3.21) into (3.19), we have
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\leq$ $(1+Ch)\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]+Ch\mathbb{E}\left[|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}|^{2}\right]+Ch^{2}$
$+C\left(\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+h\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right]\right).$
and hence for small enough $h$, we have
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\leq$ $(1+Ch)\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]+Ch^{2}$ (3.22)
$+C\left(\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+h\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right]\right).$
By Young inequality in the form: $(a+b)^{2}\geq(1-h)a^{2}+(1-\frac{1}{h})b^{2}\geq(1-h)a^{2}-\frac{1}{h}b^{2},$ for $a,b\in\mathbb{R}$, we obtain
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\geq$ $(1-h)\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]-\frac{1}{h}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right].$ (3.23)
Inserting (3.23) into (3.22), we have
$\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]$ $\leq(1+Ch)\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right|^{2}\right]+Ch^{2}$ (3.24)
$\quad+C\left(\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+h\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right]+\frac{1}{h}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\right).$
With the help of the discrete Gronwall’s lemma and (3.18), we derive
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]$ $\leq C\mathbb{E}\left[\left|\Phi(X_{T})-\Phi(X_{N}^{\pi,m})\right|^{2}\right]+Ch$ (3.25)
$\quad+C\left(\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+\frac{1}{h}\sum\limits_{i=0}^{N-1}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\right).$
Step 2. By Young inequality, we get
$\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]\leq$ $2\mathbb{E}\left[\left|\mathcal{Y}_{i}-S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)\right|^{2}\right]+2\mathbb{E}\left[\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right].$ (3.26)
In what follows, we handle each term in (3.26) separately.
[…]
$\mathbb{E}\left[\left|\mathcal{Y}_{i}-S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)\right|^{2}\right]=\mathbb{E}\left[\left|R_{i,2}^{y}\right|^{2}\right].$ (3.27)
[…]
$\mathbb{E}\left[\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]$
$\leq(1+Ch)\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta_{y,i+1}^{*})-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})-\mathbb{E}_{i}\left[\mathcal{V}^{m}_{t_{i}}\right]\right|^{2}\right]$
$\leq(1+Ch)\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|Z_{i}^{\pi,m}-R_{i,2}^{z}-R_{i,1}^{z}-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right]$
$\leq(1+Ch)\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|R_{i,1}^{z}\right|^{2}\right]+Ch\mathbb{E}\left[\left|R_{i,2}^{z}\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]\right|^{2}\right]$
$\leq(1+Ch)\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+Ch\mathbb{E}\left[\left|R_{i,1}^{z}\right|^{2}\right]+Ch\mathbb{E}\left[\left|R_{i,2}^{z}\right|^{2}\right]+C\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right|^{2}\right]$
$\leq C(1+h)\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})\right|^{2}\right]$
$\quad+C\mathbb{E}\left[\left|R_{i+1,1}^{y}\right|^{2}+h\left|R_{i,1}^{z}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,2}^{y}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}\right].$
From (3.8), we deduce
$\mathbb{E}\left[\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]$ (3.28)
$\leq C(1+h)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-R_{i+1,2}^{y}-R_{i+1,1}^{y}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,1}^{y}\right|^{2}+h\left|R_{i,1}^{z}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,2}^{2}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}\right]$
$\leq C(1+h)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-Y_{t_{i+1}}\right|^{2}+\left|Y_{t_{i+1}}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,1}^{y}\right|^{2}+h\left|R_{i,1}^{z}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,2}^{2}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}\right]$
$\leq C(1+h)+C\mathbb{E}\left[\left|R_{i+1,1}^{y}\right|^{2}+h\left|R_{i,1}^{z}\right|^{2}\right]+C\mathbb{E}\left[\left|R_{i+1,2}^{2}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}\right],$
where we use the estimates $\mathbb{E}\left[\left|Y_{i}^{\pi,m}-Y_{t_{i}}\right|^{2}\right]\leq Ch$ and $\mathbb{E}\left[\max\limits_{0\leq i\leq N}\left|Y_{t_{i}}\right|^{2}\right]<+\infty$ (see [19, 36, 28]) in the last inequalities. Plugging (3.27), (3.28) into (3.26), we derive
$\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}_{t_{i}}\right|^{2}\right]\leq$ $C\mathbb{E}\left[\left|R_{i,2}^{y}\right|^{2}+\left|R_{i+1,2}^{y}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}+\left|R_{i+1,1}^{y}\right|^{2}+h\left|R_{i,1}^{z}\right|^{2}\right].$ (3.29)
Inserting (3.29) into (3.25), we deduce
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]$ $\leq C\mathbb{E}\left[\left|g(X_{T})-g(X_{N}^{\pi,m})\right|^{2}\right]+Ch+C\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]$ (3.30)
$\quad+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|R_{i+1,2}^{y}\right|^{2}+\left|R_{i,2}^{z}\right|^{2}+N\left|R_{i+1,1}^{y}\right|^{2}+\left|R_{i,1}^{z}\right|^{2}\right].$
Step 3. Now, we prove the consistency of the $Z$-component. By (3.17) and (3.21), we obtain
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]$ (3.31)
$\leq\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+2dh\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right]$
$\quad+2d\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right|^{2}\right]$
$\quad-2d\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\leq\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+2dh\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right]$
$\quad+4d\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right|^{2}\right]+4d\mathbb{E}\left[\left|\mathcal{Y}_{i+1}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right|^{2}\right]$
$\quad-4d\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right]\right|^{2}\right]-4d\mathbb{E}\left[\left|\mathbb{E}_{i}\left[\mathcal{Y}_{i+1}-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\leq\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+2dh\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}f_{s}^{2}ds\right]$
$\quad+4d\mathbb{E}\left[\left|Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right|^{2}\right]-4d\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right]\right|^{2}\right].$
Summing over $i=0,1,\cdots,N-1$, we have
$\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]$ $\leq\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+Ch+4d\mathbb{E}\left[\left|\Phi(X_{T})-\Phi(X_{N}^{\pi,m})\right|^{2}\right]$ (3.32)
$\quad+4d\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]-4d\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right]\right|^{2}\right].$
From (3.19) and (3.23), we deduce
$4d\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]-4d\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right]\right|^{2}\right]$ (3.33)
$\leq 4d\left(\frac{1+\gamma h}{1-h}\left(1+\frac{4L^{2}h}{\gamma}\right)-1\right)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{t_{i+1}}-\mathcal{Y}_{i+1}\right]\right|^{2}\right]$
$\quad+\frac{16dL^{2}}{\gamma}\frac{1+\gamma h}{1-h}\left(Ch^{2}+\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]\right)+\frac{4d}{(1-h)h}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right].$
Taking $\gamma=48dL^{2}$, we obtain $\frac{16dL^{2}}{\gamma}\frac{1+\gamma h}{1-h}\leq\frac{1}{2}$ for sufficiently small $h$. Then plugging (3.33) into (3.32), we have
$\frac{1}{2}\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]$ (3.34)
$\leq\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]+Ch+C\max\limits_{0\leq i\leq N}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]+CN\sum\limits_{i=0}^{N-1}\mathbb{E}\left[\left|\mathcal{Y}_{i}-\mathcal{U}^{m}_{t_{i}}\right|^{2}\right]$
$\leq C\mathbb{E}\left[\left|g(X_{T})-g(X_{N}^{\pi,m})\right|^{2}\right]+Ch+C\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]$
$\quad+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|R_{i+1,2}^{y}\right|^{2}+\left|R_{i,2}^{z}\right|^{2}+N\left|R_{i+1,1}^{y}\right|^{2}+\left|R_{i,1}^{z}\right|^{2}\right],$
where we use (3.29) and (3.30) in the last inequality. From Young inequality, we know
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq 2\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]+2h\mathbb{E}\left[\left|\mathcal{V}^{m}_{t_{i}}-\mathcal{Z}_{i}\right|^{2}\right].$ (3.35)
Step 4. By Young inequality, we get
$\mathbb{E}\left[\left|\mathcal{Z}_{i}-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right]\leq$ $2\mathbb{E}\left[\left|\mathcal{Z}_{i}-S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)\right|^{2}\right]+2\mathbb{E}\left[\left|S_{z,i}\left(p^{m}_{i},\Delta W_{i}^{m},X_{i+1}^{\pi,^{m}}\right)-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right].$ (3.36)
In what follows, we handle each term in (3.36) separately.
[…]
$\mathbb{E}\left[\left|\mathcal{Z}_{i}-S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)\right|^{2}\right]=\mathbb{E}\left[\left|R_{i,2}^{z}\right|^{2}\right].$ (3.37)
[…]
$\mathbb{E}\left[\left|S_{z,i}\left(p^{m}_{i},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right]$
$=\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})\frac{(\Delta W_{i}^{m})^{\top}}{h}-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]\right|^{2}\right]$
$\leq\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right|^{2}\right]$
$\leq Nd\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})\right|^{2}\right]$
$\leq CN+CN\mathbb{E}\left[\left|R_{i+1,2}^{y}\right|^{2}+\left|R_{i+1,1}^{y}\right|^{2}\right],$ (3.38)
where we used (3.28) in the last inequality. Plugging (3.37), (3.38) into (3.36), we derive
$\mathbb{E}\left[\left|\mathcal{Z}_{i}-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right]\leq$ $CN+CN\mathbb{E}\left[\left|R_{i+1,2}^{y}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}+\left|R_{i+1,1}^{y}\right|^{2}\right].$ (3.39)
Inserting (3.39) into (3.35), we deduce
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq 2\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]+C\mathbb{E}\left[\left|R_{i+1,2}^{y}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}+\left|R_{i+1,1}^{y}\right|^{2}\right].$ (3.40)
Summing over $i=0,1,\cdots,N-1$, and then inserting (3.34) into the derived inequality, we obtain
$\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq$ $C\mathbb{E}\left[\left|g(X_{T})-g(X_{N}^{\pi,m})\right|^{2}\right]+Ch+C\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]$ (3.41)
$+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|R_{i+1,2}^{y}\right|^{2}+\left|R_{i,2}^{z}\right|^{2}+N\left|R_{i+1,1}^{y}\right|^{2}+\left|R_{i,1}^{z}\right|^{2}\right].$
Adding (3.41) to (3.30), we derive the required error estimate and this ends the proof.
3.3 Convergence analysis
In this part, we aim to study the convergence rates of the approximate error of the proposed DBR method with the help of the GroupSort deep neural networks (see [1, 17]). Now, let us briefly review it.
Let the set of the GroupSort neural networks is defined as
$\mathcal{G}_{\mathcal{K},\mathbf{d}_{0},\mathbf{d}}^{\zeta_{\kappa},L,\mathbf{n}}$= $\left\{\Psi=\left(\Psi_{i}\right)_{i=1,2,\ldots,\mathbf{d}}:\mathbb{R}^{\mathbf{d}_{0}}\mapsto\mathbb{R}^{\mathbf{d}},\Psi_{i}(x)=\mathcal{K}\overline{\beta}_{i}\psi_{i}\left(\frac{x+\alpha_{i}}{\overline{\beta}_{i}}\right),\right$.
$\left.\quad\psi_{i}\in\mathcal{S}_{\mathbf{d}_{0}}^{\zeta_{\kappa},L,\mathbf{n}}\text{ for some }\alpha_{i}\in\mathbb{R}^{\mathbf{d}_{0}},\overline{\beta}_{i}>0\right\}.$
with $\kappa\in\mathbb{N}^{*},\kappa\geq 2$, be a grouping size, dividing the number of neurons $n_{\ell}=\kappa\tilde{n}_{\ell}$, at each layer $\ell=0,\ldots,L-1$; $\sum\limits_{\ell=0}^{L-1}n_{\ell}$ denotes the width of the network and $L+1$ is the depth; $\zeta_{\kappa}=\left(\zeta_{\kappa}^{\ell}\right)_{\ell=0,\ldots,L-1}$ denotes a specific sequence of activation functions; each nonlinear function $\zeta_{\kappa}^{\ell}$ divides its input into groups of size $\kappa$ and sorts each group in decreasing order; furthermore, by enforcing the parameters of the networks to satisfy, with the Euclidian norm $|\cdot|_{2}$ and the $L_{\infty}$ norm $|\cdot|_{\infty}$,
$\mathcal{S}_{\mathbf{d}_{0}}^{\zeta_{\kappa},L,\mathbf{n}}=$ $\left\{\varphi\left(\mathcal{W}_{0},\beta_{0},\ldots,\mathcal{W}_{L},\beta_{L}\right)\in\mathcal{NN}_{\mathbf{d}_{0},1}^{\zeta_{\kappa},L,\mathbf{n}},\sup_{|x|_{2}=1}\left|\mathcal{W}_{0}x\right|_{\infty}\leq 1,\sup_{|x|_{\infty}=1}\left|\mathcal{W}_{i}x\right|_{\infty}\leq 1\right$.
$\left.\quad\left|\beta_{j}\right|_{\infty}\leq M,i=1,\ldots,L,j=0,\ldots,L\right\}.$
In what follows, we show the quantitative approximation result which is vital to study the convergence rates of the proposed deep probabilistic numerical method.
Lemma 3.2
(see Proposition 2.1 in [17]) Let $\tilde{f}:[-R,R]^{d}\rightarrow\mathbb{R}^{d_{1}}$ be $\mathcal{K}$-Lipschitz. Then, $\forall\varepsilon>0$, there exists a GroupSort neural network $g$ in $\mathcal{G}_{K,d,d_{1}}^{\zeta_{\kappa},L,\mathbf{n}}$ satisfying
$\sup\limits_{x\in[-R,R]^{d}}|\tilde{f}(x)-g(x)|\leq 2\sqrt{d_{1}}R\mathcal{K}\varepsilon,$
with $g$ of grouping size $\kappa=\lceil\frac{2\sqrt{d}}{\varepsilon}\rceil$, depth $L+1=O(d^{2})$ and width $\sum\limits_{\ell=0}^{L-1}n_{\ell}=O((\frac{2\sqrt{d}}{\varepsilon})^{d^{2}-1})$ in the case $d>1$. If $d=1$, the same result holds with $g$ of grouping size $\kappa=\lceil\frac{1}{\varepsilon}\rceil$, depth $L+1=3$ and width $\sum\limits_{\ell=0}^{L-1}n_{\ell}=O(\frac{1}{\varepsilon})$; the notation $\lceil x\rceil$ denotes the unique integer $n$ satisfying the constraint $x\leq n<x+1$.
Theorem 3.3
Suppose the assumptions (i)-(iv) hold. Furthermore, assume that $X_{0}\in L^{2+\delta}\left(\mathcal{F}_{0},\mathbb{R}^{d}\right)$, for some $\delta>0$, and $g$ is a $L_{g}$-Lipschitz function. Then, there exists a bounded sequence $\mathcal{K}_{i}$ (uniformly in $i,N$) such that for $\mathcal{NN}_{y,i}=\mathcal{G}_{\mathcal{K}_{i},1+d,n}^{\zeta_{\kappa},L,\mathbf{n}}$ and $\mathcal{NN}_{z,i}=\mathcal{G}^{\zeta_{\kappa},L,\mathbf{n}}_{\sqrt{\frac{d}{h}}\mathcal{K}_{i},1+d,n\times d}$, we have
$\max\limits_{0\leq i\leq N}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]+\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq Ch,$
with a grouping size $\kappa=O\left(2\sqrt{d}N^{3}\right)$, depth $L+1=O\left(d^{2}\right)$, and width $\sum_{\ell=0}^{L-1}n_{\ell}=$ $O\left(\left(2\sqrt{d}N^{3}\right)^{d^{2}-1}\right)$ in the case $d>1$; if $d=1$, take $\kappa=O\left(N^{3}\right)$, depth $L+1=3$, and width $\sum_{\ell=0}^{L-1}n_{\ell}=O\left(N^{3}\right)$; $K=O(N^{3})$, $K=\max\limits_{0\leq i\leq N}\{K_{i}\}$.
Proof. Step 1. Define
$\left\{\begin{aligned} \tilde{\mathcal{U}}^{m}_{t_{i}}&:=\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y+1,i})+hf\left(t_{i},\tilde{X}_{i}^{\pi,m},\tilde{\mathcal{U}}^{m}_{t_{i}},\tilde{\mathcal{V}}^{m}_{t_{i}}\right)\right],\\ \tilde{\mathcal{V}}^{m}_{t_{i}}&:=\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y+1,i})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right],\end{aligned}\right$. (3.42)
where $\tilde{p}_{i}^{m}=(t_{i},\tilde{X}_{i}^{\pi,m})$. From [31, 33, 32], we have that there exist measurable deterministic functions $y_{i}:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ and $z_{i}:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n\times d}$, such that $\tilde{\mathcal{U}}^{m}_{t_{i}}=y_{i}(t_{i},\tilde{X}_{i}^{\pi,m})$ and $\tilde{\mathcal{V}}^{m}_{t_{i}}=z_{i}(t_{i},\tilde{X}_{i}^{\pi,m})$ almost surely. From (3.10) and (3.42), we have
$h(\mathcal{V}^{m}_{t_{i}}-\tilde{\mathcal{V}}^{m}_{t_{i}})=$ $\mathbb{E}_{i}\bigg[\Big(\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})$
$-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\right]\Big)(\Delta W_{i}^{m})^{\top}\bigg],$
because the Brownian increment $\Delta W_{i}^{m}$ is conditionally centered. From Cauchy-Schwarz inequality, we have the estimate
$h|\mathcal{V}^{m}_{t_{i}}-\tilde{\mathcal{V}}^{m}_{t_{i}}|^{2}\leq$ $d\mathbb{E}_{i}\bigg[\Big(\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\Big)^{2}\bigg]$ (3.43)
$-d\bigg(\mathbb{E}_{i}\Big[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\Big]\bigg)^{2}.$
From Young inequality, we have
$|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}$ $\leq(1+\gamma h)\left|\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}$ (3.44)
$\quad+h^{2}(1+\frac{1}{\gamma h})\bigg|\mathbb{E}_{i}\big[f\left(t_{i},X_{i}^{\pi,m},\mathcal{U}^{m}_{t_{i}},\mathcal{V}^{m}_{t_{i}}\right)-f\left(t_{i},\tilde{X}_{i}^{\pi,m},\tilde{\mathcal{U}}^{m}_{t_{i}},\tilde{\mathcal{V}}^{m}_{t_{i}}\right)\Big]\bigg|^{2}$
$\leq(1+\gamma h)\left|\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}$
$+3L^{2}h^{2}(1+\frac{1}{\gamma h})\mathbb{E}_{i}\Big[|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|^{2}+|\mathcal{V}^{m}_{t_{i}}-\tilde{\mathcal{V}}^{m}_{t_{i}}|^{2}+|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}\Big].$
Plugging (3.43) into (3.44), we derive
$|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}$ $\leq(1+\gamma h-3dL^{2}(h+\frac{1}{\gamma}))\left|\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}$
$\quad+3dL^{2}(h+\frac{1}{\gamma})\mathbb{E}_{i}\left[|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})|^{2}\right]$
$\quad+3L^{2}h^{2}(1+\frac{1}{\gamma h})\mathbb{E}_{i}\left[|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|^{2}+|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}\right]$
$\leq(1+\gamma h)\mathbb{E}_{i}\left[|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})|^{2}\right]$
$\quad+3L^{2}h^{2}(1+\frac{1}{\gamma h})\mathbb{E}_{i}\left[|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|^{2}+|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}\right],$
and thus for $h$ small enough:
$|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}$ $\leq(1+Ch)\mathbb{E}_{i}\left[|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})|^{2}\right]$
$\quad+Ch\mathbb{E}_{i}\left[|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|^{2}\right],$ (3.45)
From Lemma 3.2 in [17], we get, for i=N1i=N-1
$|y(t_{N-1},X_{N-1}^{\pi,m})-y(t_{N-1},\tilde{X}_{N-1}^{\pi,m})|^{2}$ $=|\mathcal{U}^{m}_{t_{N-1}}-\tilde{\mathcal{U}}^{m}_{t_{N-1}}|^{2}$ (3.46)
$\leq(1+Ch)L_{g}^{2}|X_{N-1}^{\pi,m}-\tilde{X}_{N-1}^{\pi,m}|^{2}.$
(3.46) shows that $y(t_{N-1},X_{N-1}^{\pi,m})$ is $(1+Ch)L_{g}^{2}$-Lipschitz with respect to $x$. Similarly, if $\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y+1,i})$ is Lipschitz with respect to $x$, then $y(t_{i},X_{i}^{\pi,m})$ is also Lipschitz with respect to $x$.
Step 2. We approximate $y(t_{i},X_{i}^{\pi,m})$ by means of a $\mathcal{K}_{i}$-Lipschitz GroupSort neural network $\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i})$ in $\mathcal{NN}_{y,i}=\mathcal{G}_{\mathcal{K}_{i},1+d,n}^{\zeta_{\kappa},L,\mathbf{n}}$. From Lemma 3.2, we have the uniform error $2\mathcal{K}_{i}R\epsilon$ on $R,R]^{d}$. Then by the reverse discrete Gronwall inequality, one has
$|y(t_{i},X_{i}^{\pi,m})-y(t_{i},\tilde{X}_{i}^{\pi,m})|$ $\leq\mathcal{K}|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|,$
$|\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i})-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i},\theta^{*}_{y,i})|$ $\leq\mathcal{K}|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|,$
uniformly with respect to $i$ and $N$ for some constants $\mathcal{K}$. In the same way, we approximate $z(t_{i},X_{i}^{\pi,m})$ via a $\sqrt{\frac{d}{h}}\mathcal{K}_{i}$-Lipschitz GroupSort neural network $\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{z,i})$ in $\mathcal{NN}_{z,i}=\mathcal{G}_{\sqrt{\frac{d}{h}}\mathcal{K}_{i},1+d,n\times d}^{\zeta_{\kappa},L,\mathbf{n}}$. By Lemma 3.2, we have also the uniform error $2\frac{d}{\sqrt{h}}\mathcal{K}_{i}R\epsilon$ on $R,R]^{d}$. Therefore, $\sqrt{h}z(t,x)$ and $\sqrt{h}\mathcal{NN}_{z,t,1+d,n\times d}^{\varrho,L,\mathbf{n}}(t,x,\theta^{*}_{z})$ are $d\mathcal{K}$-Lipschitz with respect to $x$ and uniformly with respect to $N$.
Step 3. From the Hölder inequality, the Markov inequality and Lemma 3.2, we have the estimates
$\inf_{\mathcal{NN}_{y,i}\in\mathcal{G}_{\mathcal{K}_{i},1+d,n}}\mathbb{E}\left[\left|y(p^{m}_{i})-\mathcal{NN}_{y,i}(p^{m}_{i},\theta_{y,i})\right|^{2}\right]$
[…]
for $q>0$ and $2q=2+\delta$ withe $\delta$ as in the statement of the proposition. From the Lipschitz continuity of $y(t,x)$ and $\mathcal{NN}_{y,t,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{t},\theta^{*}_{y,t})$ with respect to $x$, we derive
[…]
From [30], we have the standard estimate
[…]
Hence, we deduce by (3.3)
$\inf_{\mathcal{NN}_{y,i}\in\mathcal{G}_{\mathcal{K}_{i},1+d,n}}\mathbb{E}\left[\left|y(p^{m}_{i})-\mathcal{NN}_{y,i}(p^{m}_{i},\theta_{y,i})\right|^{2}\right]\leq C\left(R^{2}\varepsilon^{2}+\frac{1+R^{2}\varepsilon^{2}}{R^{2}}\right).$ (3.49)
In the same way, repeating (3.3) and (3.48) by replacing $\mathcal{NN}_{y,i,1+d,n}(p_{i},\theta_{y,i}^{*})$ by $\mathcal{NN}_{z,i,1+d,n\times d}(p_{i},\theta^{*}_{z,i})$ and $y(t_{i},X_{i}^{\pi})$ by $z(t_{i},X_{i}^{\pi})$ and recalling that $\sqrt{h}z(t_{i},X_{i}^{\pi})$ and $\sqrt{h}\mathcal{NN}_{z,i,1+d,n\times d}(p_{i},\theta^{*}_{z,i})$ are $\mathcal{K}d$-Lipschitz with respect to $x$ and uniformly with respect to $N$, we have
$\inf_{\mathcal{NN}_{z,i}\in\mathcal{G}_{\sqrt{\frac{d}{h}}\mathcal{K}_{i},1+d,n\times d}}\mathbb{E}\left[h\left|z(p^{m}_{i})-\mathcal{NN}_{z,i}(p^{m}_{i},\theta_{z,i})\right|^{2}\right]\leq C\left(R^{2}\varepsilon^{2}+\frac{1+R^{2}\varepsilon^{2}}{R^{2}}\right).$ (3.50)
(3.49) and (3.50) show that to achieve a convergence rate of $O(\frac{1}{N})$ in Theorem 3.1, it suffices to choose the parameters $R,\varepsilon,K$ such that
$N^{2}R^{2}\varepsilon^{2}=O(\frac{1}{N}),\quad N^{2}\frac{1+R^{2}\varepsilon^{2}}{R^{2}}=O(\frac{1}{N}),\quad\frac{N^{2}}{K}=O(\frac{1}{N}),$ (3.51)
which is verified with $R=O(N^{\frac{3}{2}}),\varepsilon=O(N^{-3}),K=O(N^{3})$. From Lemma 3.2, we have that if $d>1$, we can choose the GroupSort neural network with a grouping size $\kappa=O\left(\lceil 2\sqrt{d}N^{3}\rceil\right)$, depth $L+1=O\left(d^{2}\right)$, and width $\sum_{\ell=0}^{L-1}n_{\ell}=$ $O\left(\left(2\sqrt{d}N^{3}\right)^{d^{2}-1}\right)$. If $d=1$, take $\kappa=O\left(N^{3}\right)$, depth $L+1=3$, and width $\sum_{\ell=0}^{L-1}n_{\ell}=O\left(N^{3}\right)$. From $\Phi$ satisfies the Lipschitz condition, (3.14), we obtain (see [19, 36])
$\mathbb{E}\left[\left|\Phi(X_{T})-\Phi(X_{N}^{\pi,m})\right|^{2}\right]+\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]\leq Ch.$ (3.52)
From (3.49), (3.50), (3.51) and (3.52), one yields
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]+\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq Ch,$
which completes the proof.
4 Extension to variational inequalities
In this section, we extend the DBR scheme to variational inequalities, referring to the resulting method as the RDBR scheme. Then we study the convergence of the RDBR scheme for the variational inequalities.
4.1 RDBR Scheme
Consider a variational inequality of the following form:
$\begin{cases}\min\left(-\partial_{t}u-\mu\cdot D_{x}u-\frac{1}{2}\operatorname{Tr}\left(\sigma\sigma^{\top}D_{x}^{2}u\right)+f\left(\cdot,\cdot,u,\sigma^{\top}D_{x}u\right),u-g\right)=0,&\text{ on }[0,T)\times\mathbb{R}^{d},\\ u(T,x)=g(x),&\text{ on }R^{d}.\end{cases}$ (4.1)
This inequality arises, for instance, in optimal stopping problems and American option pricing. It is established that such a variational inequality is associated with a reflected backward stochastic differential equation (RBSDE) given by
$\begin{cases}X_{t}=x_{0}+\int_{0}^{t}\mu(s,X_{s})ds+\int_{0}^{t}\sigma(s,X_{s})dW_{s},\\ Y_{t}=g(X_{T})+\int_{t}^{T}f(s,X_{s},Y_{s},Z_{s})ds-\int_{t}^{T}Z_{s}dW_{s}+K_{T}-K_{t},\\ Y_{t}\geq g(X_{t}),\quad\forall t\in[0,T],\\ \int_{0}^{T}(Y_{t}-g(X_{t}))dK_{t}=0,\end{cases}$ (4.2)
where $K$ is an adapted non-decreasing process satisfying the Skorokhod condition (see [14]). The Euler time-discretization of the RFBSDE (4.2), at the mesh points $t_{i}$, is
$Y_{i}^{\pi}=Y_{i+1}^{\pi}+hf(t_{i},X_{i}^{\pi},Y_{i}^{\pi},Z_{i}^{\pi})-Z_{i}^{\pi}\Delta W_{i}+K_{i+1}^{\pi}-K_{i}^{\pi},$
which also reads as the conditional expectation formula
$\left\{\begin{aligned} \overline{Y}_{i}^{\pi}=&\mathbb{E}_{i}\left[Y^{\pi}_{i+1}+hf(t_{i},X_{i}^{\pi},\overline{Y}_{i}^{\pi},Z_{i}^{\pi})\right],\\ Y_{i}^{\pi}=&\max\left\{\overline{Y}^{\pi}_{i},g\left(X_{i}^{\pi}\right)\right\},\\ Z_{i}^{\pi}=&\mathbb{E}_{i}\left[Y_{i+1}^{\pi}\frac{\Delta W_{i}^{\top}}{h}\right].\end{aligned}\right$.
The backward process $Y_{\cdot}^{\pi,m}$ in (4.2) is simulated by Monte Carlo method in the following ways, for $i=N-1,N-2,\cdots,0$
$Y_{i}^{\pi,m}=Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})-Z_{i}^{\pi,m}\Delta W_{i}^{m}+K_{i+1}^{\pi,m}-K_{i}^{\pi,m},$
which also expresses as the conditional expectation form
$\left\{\begin{aligned} \overline{Y}_{i}^{\pi,m}=&\mathbb{E}_{i}\Big[Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m})\Big],\\ Y_{i}^{\pi,m}=&\max\left\{\overline{Y}^{\pi,m}_{i},g\left(X_{i}^{\pi,m}\right)\right\},\\ Z_{i}^{\pi,m}=&\mathbb{E}_{i}\big[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\big].\end{aligned}\right$. (4.3)
Thus, we derive the fully discrete approximation $(\widehat{\mathcal{Y}}_{i},\widehat{\mathcal{Z}}_{i})$ for the RDBR scheme for the solutions $(Y_{t_{i}},Z_{t_{i}})$ of the variational inequality (4.1) as follows, for $i=N,N-1,\cdots,1,0$, $m=1,2,\cdots,1,M$
  1. 1.
    the terminal condition is $(\widehat{\mathcal{Y}}_{N},\widehat{\mathcal{Z}}_{N})=(\Phi(X_{N}^{\pi,m}),\sigma^{\top}(t_{N},X^{\pi,m}_{N})D_{x}u(t_{N},X^{\pi,m}_{N}))$,
  2. 2.
    for $0\leq i<N,$ the transition from $i+1$ to $i$ is given by
    $\left\{\begin{aligned} \widehat{\mathcal{Y}}_{i}=&\max\left\{\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i}),g\left(X_{i}^{\pi,m}\right)\right\},\\ \widehat{\mathcal{Z}}_{i}=&\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{z,i}).\end{aligned}\right$. (4.4)
    and compute (by SGD) the minimizer of the quadratic loss function, for $r_{j}\in\mathbb{N}^{+}$ large enough
    $\left\{\begin{aligned} \widehat{F}_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})=&\frac{1}{M}\sum_{m=1}^{M}\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2},\\ \widehat{F}_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})=&\frac{1}{M}\sum_{m=1}^{M}\left|S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)-\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2},\end{aligned}\right$.
    where $\theta^{*}_{y,i}\in\arg\min\limits_{\theta_{y,i}^{r_{j}}}\widehat{F}_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})$, $\theta^{*}_{z,i}\in\arg\min\limits_{\theta_{z,i}^{r_{j}}}\widehat{F}_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})$.
4.2 Convergence analysis
In this section, we analyze the convergence of the RDBR schemes (4.4) for the variational inequality (4.1) related to the solution $(Y,Z)$ of the RFBSDE (4.2).
Now, we investigate the errors of the scheme (4.4) and define the auxiliary quantities as below, for $i=N-1,N-2,\cdots,0$
$\left\{\begin{aligned} \overline{\mathcal{U}}_{t_{i}}^{m}&=\mathbb{E}_{i}\left[\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})+hf\left(t_{i},X_{i}^{\pi,m},\overline{\mathcal{U}}_{t_{i}}^{m},\widehat{\mathcal{V}}_{t_{i}}^{m}\right)\right],\\ \widehat{\mathcal{U}}_{t_{i}}^{m}&=\max\left\{\overline{\mathcal{U}}_{t_{i}}^{m},g\left(X_{i}^{\pi,m}\right)\right\},\\ \widehat{\mathcal{V}}_{t_{i}}^{m}&=\mathbb{E}_{i}\left[\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right].\end{aligned}\right$. (4.5)
The result is obtained under one of the following additional assumptions:
Assumption 4.1
$g$ is $C^{1}$, and $g,D_{x}g$ are Lipschitz.
Assumption 4.2
$\sigma$ is $C^{1}$, with $\sigma,D_{x}\sigma$ both Lipschitz, and $g$ is $C^{2}$, with $g,D_{x}g,D^{2}_{x}g$ all Lipschitz.
Theorem 4.3
Let the assumptions (i)-(iv) hold. Let $(Y_{t_{i}},Z_{t_{i}})$ and $(\widehat{\mathcal{Y}}_{i},\widehat{\mathcal{Z}}_{i})$ be solutions of the RFBSDE (4.2) and solutions of the RDBR method (4.4) respectively. Then, we have, for small enough $h$
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]+\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\widehat{\mathcal{Z}}_{i}|^{2}ds\right]$
$\leq C\varepsilon(\pi)+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|\widehat{R}_{i+1,2}^{y}\right|^{2}+\left|\widehat{R}_{i,2}^{z}\right|^{2}+N\left|\widehat{R}_{i+1,1}^{y}\right|^{2}+\left|\widehat{R}_{i,1}^{z}\right|^{2}\right],$
with $\varepsilon(h)=O(h^{1/2})$ under Assumption 4.1, and $\varepsilon(h)=O(h)$ under Assumption 4.2;
$\widehat{R}_{i,1}^{y}$ $=\mathbb{E}_{i}\Big[Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m})\Big]-\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}-hf(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m}),$
$\widehat{R}_{i,1}^{z}$ $=\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]-\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right],$
$\widehat{R}_{i,2}^{y}$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}+hf(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m})-\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{*}),$
$\widehat{R}_{i,2}^{z}$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right]-\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*}).$
Proof. From [3], we have
$\begin{cases}\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-Y_{i}^{\pi,m}\right|^{2}\right]&=\varepsilon(h),\\ \mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{t}-Z^{\pi,m}_{i}|^{2}dt\right]&=O(h^{\frac{1}{2}}),\end{cases}$ (4.6)
with $\varepsilon(h)=O(h^{1/2})$ under Assumption 4.1, and $\varepsilon(h)=O(h)$ under Assumption 4.2. From (4.3) and (4.5), we have, for $i\in\{0,1,\ldots,N-1\}$
$\overline{Y}_{i}^{\pi,m}-\overline{\mathcal{U}}^{m}_{t_{i}}=$ $\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]$ (4.7)
$+h\left(f(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m})-f\left(t_{i},X_{i}^{\pi,m},\overline{\mathcal{U}}_{t_{i}}^{m},\widehat{\mathcal{V}}_{t_{i}}^{m}\right)\right).$
Proceeding similarly as the Step 1 in the proof pf Theorem 3.1, we obtain by Young inequality, Cauchy-Schwarz inequality, and the Lipschitz condition with respect to $f$
$\mathbb{E}\left[\left|\overline{Y}_{i}^{\pi,m}-\overline{\mathcal{U}}^{m}_{t_{i}}\right|^{2}\right]$ $\leq(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$
$\quad+2L^{2}h^{2}(1+\frac{1}{\gamma h})\mathbb{E}\left[|\overline{Y}_{i}^{\pi,m}-\overline{\mathcal{U}}^{m}_{t_{i}}|^{2}+|Z_{i}^{\pi,m}-\widehat{\mathcal{V}}^{m}_{t_{i}}|^{2}\right].$ (4.8)
From Cauchy-Schwarz inequality and the tower property of conditional expectations, one yields
$\mathbb{E}\left[\left|Z_{i}^{\pi,m}-\widehat{\mathcal{V}}^{m}_{t_{i}}\right|^{2}\right]\leq$ $\frac{2d}{h}\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]$ (4.9)
$-\frac{2d}{h}\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right].$
Plugging (4.9) into (4.8) and choosing $\gamma=4dL^{2}$, we have for small enough $h$
$\mathbb{E}\left[\left|\overline{Y}_{i}^{\pi,m}-\overline{\mathcal{U}}^{m}_{t_{i}}\right|^{2}\right]\leq$ $(1+Ch)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right].$ (4.10)
Next, by Young inequality as the Step 1 in the proof of Theorem 3.1, we derive
$\mathbb{E}\left[\left|\overline{Y}_{i}^{\pi,m}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]$ $\leq(1+Ch)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]$ (4.11)
$\quad+CN\mathbb{E}\left[\left|\widehat{\mathcal{Y}}_{i}-\overline{\mathcal{U}}^{m}_{t_{i}}\right|^{2}\right].$
By the same arguments as in Step 2 in the proof of Theorem 3.1, it follows that for $h$ small enough
$\mathbb{E}\left[\left|\widehat{\mathcal{Y}}_{i}-\overline{\mathcal{U}}^{m}_{t_{i}}\right|^{2}\right]$ $\leq C\mathbb{E}\left[N\left|\widehat{R}_{i+1,2}^{y}\right|^{2}+\left|\widehat{R}_{i,2}^{z}\right|^{2}+N\left|\widehat{R}_{i+1,1}^{y}\right|^{2}+\left|\widehat{R}_{i,1}^{z}\right|^{2}\right].$ (4.12)
Plugging (4.12) into (4.11) and recalling that $\widehat{\mathcal{U}}_{t_{i}}^{m}=\max\left\{\overline{\mathcal{U}}_{t_{i}}^{m},g\left(X_{i}^{\pi,m}\right)\right\}$, $Y_{i}^{\pi,m}=\max\left\{\overline{Y}^{\pi,m}_{i},g\left(X_{i}^{\pi,m}\right)\right\}$ and $\left|\max(a,c)-\max(b,c)\right|\leq\left|a-b\right|$, we have
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{i}^{\pi,m}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]\leq C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|\widehat{R}_{i+1,2}^{y}\right|^{2}+\left|\widehat{R}_{i,2}^{z}\right|^{2}+N\left|\widehat{R}_{i+1,1}^{y}\right|^{2}+\left|\widehat{R}_{i,1}^{z}\right|^{2}\right].$ (4.13)
Combining (4.13) with (4.6), we deduce
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]\leq C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|\widehat{R}_{i+1,2}^{y}\right|^{2}+\left|\widehat{R}_{i,2}^{z}\right|^{2}+N\left|\widehat{R}_{i+1,1}^{y}\right|^{2}+\left|\widehat{R}_{i,1}^{z}\right|^{2}\right].$ (4.14)
This establishes the error estimate for the $Y$-component. The error estimate for the $Z$-component can be derived using the same arguments as in Steps 3 and 4 of the proof of Theorem 3.1; thus, the details are omitted here.
5 Numerical experiments
In this section, we demonstrate the performance of Algorithm 1 to solve the high dimensional nonlinear PDEs (1.1) via two numerical examples.
All the numerical tests are implemented in Python 3.8 on a desktop computer with Intel(R) Xeon(R) Gold 6133 CPU (2.50GHz) and 64 GB RAM (3200MHz), and an NVIDIA GeForce RTX 4090 GPU (24 GB) with CUDA 11.8 support. All implementations were developed using PyTorch 2.0. The corresponding code is available at: https://github.com/22w2e/A-DNMC-algorithm-for-high-dimensional-nonlinear-PDEs and the results are recorded. Note that we still implement the proposed deep method by the feedforward neural network because the computational cost of the GroupSort network is bigger than that of the costly feedforward neural network in practice (see [17]). To mitigate randomness inherent in stochastic sampling, each experiment is independently repeated $10$ times under identical conditions, and the corresponding results are reported. Although a larger sample size would tighten the confidence intervals, the computational cost for these high-dimensional problems is significant. Ten repetitions allow us to obtain a reliable estimate of the mean error magnitude and variance, which is consistent with standard practices in related literature [28].
Example 1. We first consider the following $d$-dimensional nonlinear parabolic PDE (taken from [11]):
$\begin{cases}\displaystyle\frac{\partial u}{\partial t}(t,x)+\frac{d^{2}}{2}\,\Delta_{x}u(t,x)+\Big(u(t,x)-\frac{d+2}{2d}\Big)\Big(d\sum_{\ell=1}^{d}\frac{\partial u}{\partial x_{\ell}}(t,x)\Big)=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\[10.00002pt] \displaystyle u(T,x)=\frac{\exp\!\left(T+\tfrac{1}{d}\sum_{\ell=1}^{d}x_{\ell}\right)}{1+\exp\!\left(T+\tfrac{1}{d}\sum_{\ell=1}^{d}x_{\ell}\right)}.\end{cases}$ (5.1)
The corresponding analytic solution is explicitly given by
$u(t,x)=\frac{\exp\!\left(t+\tfrac{1}{d}\sum_{\ell=1}^{d}x_{\ell}\right)}{1+\exp\!\left(t+\tfrac{1}{d}\sum_{\ell=1}^{d}x_{\ell}\right)}.$
In the numerical experiments we take the terminal time $T=1$, the number of time steps $N=10$, the Monte Carlo samples $M=10000$, and the initial state $x_{0}=(0,0,\dots,0)^{\top}\in\mathbb{R}^{d}$. Thus, The true solution of (5.1) at the initial point is $u(0,x_{0})=0.5$. The numerical solution at the initial point is denoted by $\mathfrak{u}(0,x_{0})$. The numerical solver is implemented using a deep learning framework where two sub-networks are trained at each time step to approximate the solution $u$ and its gradient $\nabla u$. Each network consists of two hidden layers with $d+110$ units and $\tanh$ activations. We utilize the Adam optimizer with a learning rate of $5\times 10^{-4}$ and conduct 6,000 training iterations per step to ensure the convergence of the loss function.
$u(t,x)$ and its estimate at time $t=0.2$.
$u(t,x)$ and its estimate at time $t=0.4$.
$u(t,x)$ and its estimate at time $t=0.6$.
$u(t,x)$ and its estimate at time $t=0.8$.
Figure 1: Estimated solution $\mathfrak{u}(t,x)$ obtained by DBR versus exact solution $u(t,x)$ for Example 1 with $d=1$.
Figure 1 provides a qualitative visualization of the DBR algorithm’s approximation accuracy across the entire time domain. The figure displays the estimated solution $\overline{\mathfrak{u}}(t,x)$ alongside the exact analytical solution $u(t,x)$ at four distinct time snapshots ($t=0.2,0.4,0.6,0.8$) for the one-dimensional case. Across all four subplots, the estimated values exhibit near-perfect alignment with the true solution, with no visible divergence or significant error at any intermediate time step. This visual evidence demonstrates that the high accuracy achieved by the DBR method is not confined to the initial condition at $t=0$, but is consistently maintained throughout the entire time horizon from $t=0$ to $t=T$. The results confirm the DBR scheme’s capability to correctly propagate the solution backward in time through its learned network functions, validating the effectiveness of the expectation-based loss functions in capturing the underlying dynamics of the PDE.
Table 1: Comparison of numerical results for different schemes in Example 1 across different dimensions.
Scheme $d$ $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
DBR 50 0.500000 0.517697 0.007637 0.017697 3.5393%
80 0.500000 0.508568 0.007589 0.008683 1.7137%
100 0.500000 0.507220 0.005692 0.008006 1.4441%
200 0.500000 0.482896 0.007052 0.017104 3.4207%
DBDP1 [28] 50 0.500000 0.569523 0.008724 0.069523 13.9047%
80 0.500000 0.548494 0.005876 0.048494 9.6989%
100 0.500000 0.545578 0.005622 0.045578 9.1155%
200 0.500000 0.524425 0.002196 0.024425 4.8850%
Table 1 complements these visual findings by providing a quantitative comparison between the proposed DBR algorithm and the existing DBDP1 method in [28] across multiple high dimensions ($50,80,100$ and $200$). For each scheme and dimension, we report the sample mean $\overline{\mathfrak{u}}(0,x_{0})$ over 10 independent runs, the empirical standard deviation $\operatorname{Std.dev.}(u^{\pi}(0,x_{0}))$, the mean absolute error $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$, and the relative error $\operatorname{Rel.err.}(\mathfrak{u}(0,x_{0}))=|u(0,x_{0})-\overline{\mathfrak{u}}(0,x_{0})|/|u(0,x_{0})|$. Regarding accuracy, the DBR algorithm maintains a relative error consistently below $4\%$ across all tested dimensions, significantly outperforming the DBDP1 method which exhibits errors as high as $13.9\%$ at dimension $50$ and $9.7\%$ at dimension $80$. More importantly, the mean absolute error, which represents the average magnitude of the absolute deviation between the true solution and the numerical approximation, is consistently smaller for the DBR method across all dimensions. For instance, at dimension $100$, DBR achieves a mean absolute error of $0.008006$ compared to DBDP1’s $0.045578$, indicating that the DBR algorithm produces approximations that are substantially closer to the true solution on average. This aligns with the papers’ claim that the DBR method produces more accurate estimations, particularly as the dimension increases. Regarding stability, the standard deviations reported for both methods are generally comparable in magnitude across the tested dimensions. The numerical results validate the core theoretical advantages of the DBR method.
Example 2. (see [17]) Consider the following high-dimensional nonlinear PDEs which are equipped with unbounded solution and more complex structure of the form
[…]
where the generator
$f(t,x,y,z)=\hat{k}(t,x)-\frac{y}{\sqrt{d}}\left(1_{d}\cdot z\right)-\frac{y^{2}}{2},$
with $\hat{k}(t,x)=\partial_{t}u+\frac{1}{2d}\operatorname{Tr}\left(D_{x}^{2}u\right)+\frac{u}{\sqrt{d}}D_{x_{i}}u+\frac{u^{2}}{2}$. Thus, the solution of PDE (5.2) is given by $u(t,x)=\frac{T-t}{d}\sum_{i=1}^{d}\left(\sin\left(x_{i}\right)1_{x_{i}<0}+x_{i}1_{x_{i}\geq 0}\right)+\cos\left(\sum_{i=1}^{d}ix_{i}\right).$
Example 2 presents a more challenging class of high-dimensional nonlinear PDEs characterized by an unbounded solution and a complex structural form. To comprehensively evaluate the performance of the proposed DBR algorithm, we conduct systematic numerical experiments across dimensions $d=1,2,8,15,20$ with varying time steps $N$, where the initial point is fixed as $x_{0}=0.5\mathbf{1}_{d}$ for all cases. The neural network architecture consists of two hidden layers with hyperbolic tangent activation functions, where each hidden layer contains $d+10$ units. The model is trained using a batch size of 400 and a total of 3000 iterations.
$u(t,x)$ and its estimate at time $t=0.3$.
$u(t,x)$ and its estimate at time $t=0.5$.
$u(t,x)$ and its estimate at time $t=0.7$.
$u(t,x)$ and its estimate at time $t=0.9$.
Figure 2: Estimated solution $\overline{\mathfrak{u}}(t,x)$ obtained by DBR versus exact solution $u(t,x)$ for Example 2 with $d=1$.
Table 2: Comparison of numerical results for different schemes in Example 2 across different NN with $d=1$.
Scheme NN $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
DBR 10 1.377583 1.403209 0.011086 0.025700 1.8602%
20 1.377583 1.400363 0.006224 0.022780 1.6536%
30 1.377583 1.393180 0.006930 0.015598 1.1323%
DBDP1 [28] 10 1.377583 1,371497 0.023389 0.019112 0.4417%
20 1.377583 1.366828 0.017810 0.017439 0.7806%
30 1.377583 1.367642 0.019181 0.017814 0.7216%
Table 3: Comparison of numerical results for different schemes in Example 2 across different NN with $d=2$.
Scheme NN $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
DBR 10 0.570737 0.571529 0.014380 0.012455 0.1387%
20 0.570737 0.579191 0.016317 0.015263 1.4813%
30 0.5707373 0.574875 0.016678 0.014892 0.7250%
DBDP1 [28] 10 0.570737 0.560685 0.009967 0.011556 1.7613%
20 0.570737 0.563375 0.009665 0.009730 1.2899%
30 0.570737 0.567637 0.007557 0.006504 0.5431%
Table 4: Comparison of numerical results for different schemes in Example 2 across different NN with $d=8$.
Scheme NN $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
DBR 70 1.160317 1.240760 0.150684 0.131828 6.9328%
90 1.160317 1.211207 0.116388 0.092109 4.3859%
110 1.160317 1.232262 0.111441 0.113141 6.2005%
DBDP1 [28] 70 1.160317 1.126632 0.134559 0.107561 2.9030%
90 1.160317 1.173978 0.077135 0.061737 1.1773%
110 1.160317 1.121225 0.090274 0.075801 3.3690%
Figure 2 provides a qualitative visualization of the approximation accuracy achieved by the DBR algorithm for the one-dimensional case of Example 2. The figure displays the estimated solution $\overline{\mathfrak{u}}(t,x)$ obtained by the DBR method alongside the exact analytical solution $u(t,x)$ at four distinct time snapshots (t=0.3, 0.5, 0.7, 0.9). This visual evidence confirms that the DBR algorithm successfully captures the underlying dynamics of the PDE throughout the entire time horizon, demonstrating its capability to accurately propagate the solution backward in time through the learned network functions. The excellent agreement between the estimated and exact solutions serves as compelling qualitative validation of the effectiveness of the expectation-based loss functions employed in the DBR methodology.
For low-dimensional cases ($d=1,2,8$), both the DBR and DBDP1 methods demonstrate satisfactory performance, achieving relative errors generally below $7\%$ across different $N$. As shown in Tables 2, 3 and 4, the two algorithms exhibit comparable accuracy and stability, with no single method displaying a decisive advantage in these low-dimensional regimes. This observation aligns with the theoretical expectation that both probabilistic numerical schemes are well-suited for problems of low dimensionality.
Table 5: Comparison of numerical results for different schemes in Example 2 across different NN with $d=15$.
Scheme NN $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
DBR 120 -0.452413 -0.491338 0.125350 0.112779 8.6040%
130 -0.452413 -0.496170 0.194659 0.164798 9.6719%
140 -0.452413 -0.458081 0.245985 0.225215 1.2529%
DBDP1 [28] 120 -0.452413 -0.730808 0.146088 0.278396 61.5357%
130 -0.452413 -0.653904 0.207671 0.271590 44.5369%
140 -0.452413 -0.703535 0.188429 0.271442 55.5073%
Table 6: Comparison of numerical results for different schemes in Example 2 across different NN with $d=20$.
Scheme NN $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
DBR 90 0.259041 0.249463 0.446021 0.366364 3.6974%
100 0.259041 0.267776 0.417597 0.305940 3.3720%
140 0.259041 0.235651 0.264053 0.216113 9.0293%
DBDP1 [28] 90 0.259041 0.076700 0.288634 0.264827 70.3907%
100 0.259041 0.116744 0.230302 0.234747 54.9322%
140 0.259041 0.051147 0.286756 0.314811 80.2554%
However, a markedly different pattern emerges as the dimension increases beyond $10$. For $d=15$ (Table 5), the DBDP1 method exhibits a dramatic deterioration in performance, with relative errors soaring to approximately $61.5357\%$, $44.5369\%$ and $55.5073\%$ for $N=120,130,140$, respectively. These unacceptably large errors indicate that DBDP1 essentially fails to provide reliable approximations in this higher-dimensional setting, a finding that corroborates the observations in [28] where the authors noted that DBDP1 becomes ineffective when the dimension exceeds 10. In stark contrast, the DBR method maintains relative errors consistently below $10\%$ across all tested temporal discretizations, achieving an impressive $1.2529\%$ relative error at $N=140$. This remarkable stability and accuracy demonstrate the robustness and superior approximation capability of DBR in high-dimensional regimes.
The advantage of DBR becomes even more pronounced at $d=20$ (Table 6). While DBDP1 produces relative errors of $70.3907\%$, $54.9322\%$ and $80.2554\%$ for $N=90,100,140$, rendering its estimates practically meaningless, DBR achieves remarkably low relative errors of $6974\%, 3.3$$3720\%$ and $9.0293\%$ under the same conditions. This striking performance gap unequivocally demonstrates that the DBR algorithm retains its accuracy and stability in high-dimensional regimes where traditional methods like DBDP1 completely break down.
These numerical results collectively substantiate the core contribution of this work: by replacing the pathwise loss function with expectation-based alternatives, the DBR method effectively mitigates the adverse effects of increasing dimensionality, thereby offering a reliable and accurate approach for solving high-dimensional nonlinear PDEs. The method’s consistent performance across dimensions $d=15$ and $d=20$, particularly when compared to the catastrophic failure of DBDP1, provides compelling evidence of its practical utility for real-world high-dimensional problems. Due to prohibitive computational costs, the high-dimensional regime ($d>20$) was excluded from our experiments. We anticipate that the results would benefit from an increased sample size.
6 Conclusion
In this paper, we have proposed a new deep probabilistic numerical algorithm which is called deep backward regression-based method for high-dimensional quasi-linear PDEs based on the probabilistic representation of the solution of the PDEs. The computation of our loss functions essentially involves performing a local Monte Carlo simulation at each time step, using the average value to replace the individual sample value for calculating the labels. Its advantage is that we do not need to fit the Brownian motion (Since the noise is not fittable, this wastes the capacity of the algorithm.) and we merely need to fit the average dynamics of the system. Thus, our DBR method has a higher learning efficiency. Then we have rigorously proven that the error estimates of the proposed DBR method. Moreover, we have established the half-order convergence of our deep numerical scheme by rigorous analysis. Furthermore, it is interesting and challenging that one develops an optimization of sampling strategy to reduce the computational cost of conditional expectations or an lightweight the neural network architecture to reduce computational cost of the models to improve the the numeral method of PDEs.
Declarations
Conflict of interest
The authors declare that they have no conflict of interest.
Data Availability
Data sharing is not applicable to this article as no datasets are generated or analyzed in this study.
Authors’ contributions
The authors contributed equally to this article. The authors read and approved the final manuscript.
Funding
This work was supported by the National Key R&D Program of China (No.2023YFA1008701), the Key Project of the National Natural Science Foundation of China (No.12431017), the National Natural Science of China (No.12501664), the National Natural Science of Yangzhou, China (No.YZ2025145) and the Golden Phoenix of the Green City-Yang Zhou (No.137013391).